OpenStudy (anonymous):

Find the solutions of g(x). Show each step. g(x)x^2-6x=-16

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

\[x^2-6x=-16\]Add 16 to both sides to get in desired form:\[x^2-6x+16 = -16+16\]\[x^2-6x+16=0\]\[ax^2+bx+c=0\]\[a=1,b=-6,c=16\]Use the quadratic formula to get the solutions:\[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]